OpenStudy (anonymous):

medal! for the best answer A can is full of spirit. Out of this, 7 litres of spirit are replaced with water. If the process is repeated 3 times, the ratio of spirit to water becomes 27 : 37. What is the full capacity of the can?

OpenStudy (anonymous):

x(1-7/x)/x=27/37

OpenStudy (anonymous):

(1-7/x)^3=27/37

OpenStudy (anonymous):

(x-7/x)=3/37^1/3

ganeshie8 (ganeshie8):

hold up, lets work the first problem

ganeshie8 (ganeshie8):

or u figured out how to solve first problem already ? :)